Seafoam:

Use the chart below to answer the questions: Country Population (1997 estimate) Area (square km) Indonesia 203,479,000 1,904,569 Denmark 5,248,000 42,077 Ghana 18,339,000 237,533 Which of these countries has the greatest population density? Which has the least? Provide an explanation and proof for your answer.

Seafoam:

@AZ

Seafoam:

Ok I'm gonna try to rearrange the chart, that one doesnt make much sense

Seafoam:

Country Population Area (1997 estimate) (square km) Indonesia 203,479,000 1,904,569 Denmark 5,248,000 42,077 Ghana 18,339,000 237,533

AZ:

It's population and then the area in square miles so Indonesia has a population of 203 million and a land area of 1.9 million square km

AZ:

ye that does look much better now

AZ:

so population density is number of people per unit area so just divide the first number by the second number that will tell you how many people on average live in 1 square km that's the population density and then you can figure out which country has the highest population density and the lowest

Seafoam:

Indonesia is 106.8

Seafoam:

And denmark is 124.7

Seafoam:

And Ghana has the least with 77.2

Seafoam:

How do you do the division symbol?

Seafoam:

Is it ^?

AZ:

that would be exponent division is using this /

AZ:

But those three numbers look correct! So which one is the largest and which one is the least?

Seafoam:

oh, oh i forgot šŸ˜…

Seafoam:

Indonesia'sĀ population density= 203,479,000 / 1,904,569= 106.8 Denmark'sĀ population density= 5,248,000 /Ā 42,077= 124.7 Ghana'sĀ population density= 18,339,000 / 237,533= 77.2

Seafoam:

Is that all?

AZ:

yup!

Seafoam:

Well I added Ghana has the least in a sentance but other than that is there anything else?

Seafoam:

Ok great

AZ:

just add that Denmark has the highest too

Seafoam:

Haha I've seen that one first hand, Copenhagens prices are just ridiculous